OpenStudy (anonymous):

(Geometry) Enter the equation of the circle in standard form with center and radius given. Center (7, 5), radius = 4

OpenStudy (anonymous):

\[\frac{(x-7)^2}{16}+\frac{(y-5)^2}{16}=1\]

OpenStudy (anonymous):

Thank you!